Consider the following information:

  

Rate of Return if State Occurs
State of Economy Probability of
State of Economy
Stock A Stock B Stock C
Boom 0.70 0.07 0.09 0.29
Bust 0.30 0.19 0.09 0.17

  

Requirement 1:

What is the expected return on an equally weighted portfolio of these three stocks? (Do not round your intermediate calculations.)

(Click to select)15.00%17.50%27.03%29.80%9.27%

  

Requirement 2:

What is the variance of a portfolio invested 20 percent each in A and B and 60 percent in C? (Do not round your intermediate calculations.)

(Click to select)0.0039840.0084840.0059840.0004840.008184
Answer #1

a) Boom return on the portfolio = ( 7% +9% + 29% ) /3= 15%

Bust return on portfolio = ( 19% +9% +17% ) /3 = 15%

expected return on an equally weighted portfolio of these three stocks = 0.7 * 15% + 0.3 * 15% = 15%

Hi, please provide correct options for B. not able to understand